7.3 kg of copper sits at a temperature of 38 degrees F. How much heat is required to raise its temperature to 865 degrees F? The specific heat of copper is 385 J/kg- degree C. Submit your anwser in exponential form.


1
Expert's answer
2022-01-31T09:05:16-0500

Calculate the heat:


"Q=cm(T_f-T_i),\\
Q=385·7.3·(735.9-276.5)=1.29·10^6\text{ J}."